OpenStudy (mitu12):

solve the inequality 3r+3 greater than or equal to sign -12

OpenStudy (mitu12):

@Castiel

OpenStudy (mitu12):

@OOOPS

OpenStudy (anonymous):

do you have an idea on what we do first

OpenStudy (mitu12):

you subtract 3 from -12

OpenStudy (anonymous):

yes \[3r \ge-12\] now what

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

\[3r \ge-15\]

OpenStudy (mitu12):

i think the answer is 3r =-f but i dont know when to flip the sign

OpenStudy (mitu12):

-5

OpenStudy (anonymous):

we only flip the sign when we divide by a negative

OpenStudy (mitu12):

ohhh so its

OpenStudy (anonymous):

\[-3r \ge-15\] would turn into \[r \le5\]

OpenStudy (mitu12):

r greater then or equal to -5

OpenStudy (anonymous):

yes \[r \ge-5\]

OpenStudy (anonymous):

Perfect

OpenStudy (mitu12):

thank you so much

OpenStudy (anonymous):

you're welcome